A sodium atom has a mass of 22.99 u. Use Avogadro’s constant and the conversion factor
1.661 x 10–24 g/u to calculate the mass of 1.00 mol of sodium atoms.
1
Expert's answer
2020-08-11T04:35:20-0400

Let us find the actual mass of the sodium atom:


"m_{0}(Na)=22.99u*1.661*10^{-24}g\/u=3.819*10^{-23}g;"


This as a single atom, while 1 mole contains of Avogadro number of atoms:


"M_{1mol}(Na)=m_{0}(Na)*N_{A}=3.819*10^{-23}g*6.02*10^{23}=22.99g;"


Thus, the mole of sodium atoms weights 22.99 grams.